Bibliography

[1] W. Athas. Low-power VLSI techniques for applications in embedded computing. In IEEE Alessandro Volta Memorial Workshop on Low-Power Design, pages 14–22. IEEE Computer Society Press, March 1999. Abstract at http://www.-

isi.edu/acmos/abstracts/99-04.VOLTA.html.

[2] W. Athas, N. Tzartzanis, L. Svensson, L. Peterson, H. Li, X. Jiang, P. Wang, and W-C. Liu. AC-1: A clock-powered microprocessor. In Proc. of the International Symposium on Low-Power Electronics and Design, Monterey, CA, 18–

20 August 1997. http://www.isi.edu/acmos/papers/97-08.MontereyAC1.-
ps.

[3] W. C. Athas, L. “J.” Svensson, and N. Tzartzanis. A resonant signal driver for two-phase, almost-non-overlapping clocks. In ISCAS-96, 1996.

[4] W. C. Athas, N. Tzartzanis, L. “J.” Svensson, and L. Peterson. A low-power microprocessor based on resonant energy. IEEE Journal of Solid-State Circuits, 32(11):1693–1701, nov 1997.

[5] William C. Athas, Lars “J.” Svensson, Jeffrey G. Koller, Nestoras Tzartzanis, and Eric Ying-Chin Chou. Low-power digital systems based on adiabaticswitching principles. IEEE Transactions on Very Large Scale Integration (VLSI) Systems, 2(4):398–407, December 1994.

[6] Stephen Avery and Marwan Jabri. Design of a register file using adiabatic logic. Technical Report SCA-06/06/97, University of Sydney SEDAL, June 1997.

[7] Henry G. Baker. Lively linear lisp—‘Look ma, no garbage!’. ACM SigPlan Notices, 27(8):89–98, August 1992.

[8] Henry G. Baker. NREVERSAL of fortune — the thermodynamics of garbage collection. In Y. Bekkers, editor, International Workshop on Memory Management, pages 507–524. Springer-Verlag, 1992.

[9] Theodore Baker, John Gill, and Robert Solovay. Relativizations of the P =?N P question. SIAM J. Computing, 4(4):431–442, December 1975.

[10] Adriano Barenco, David Deutsch, Artur K. Ekert, and Richard Jozsa. Conditional quantum dynamics and logic gates. Physical Review Letters, 74(20):4083– 4086, 15 May 1995. Preprint at Los Alamos Physics Preprint Archive, http:/

/xxx.lanl.gov/abs/quant-ph/9503017.

[11] Edward Barton. A reversible computer using conservative logic. Term paper for 6.895 at MIT, 1978.

[12] Matthew E. Becker and Thomas F. Knight, Jr. Transmission line clock driver. In Power Driven Microarchitecture Workshop, pages 80–85, Barcelona, Spain, 28 June 1998.

[13] Matthew E. Becker and Thomas F. Knight, Jr. Transmission line clock driver. In IMAPS Int’l Advanced Technology Wkshp. on Flip Chip Technology, Braselton, GA, 12–14 March 1999.

[14] Jacob D. Bekenstein. Universal upper bound on entropy-to-energy ratio for bounded systems. Phys. Rev. D, 23(2):287–298, January 1981.

[15] Jacob D. Bekenstein. Entropy content and information flow in systems with limited energy. Phyical Review D, 30(8):1669–1679, 15 October 1984.

[16] C. H. Bennett. Logical reversibility of computation. IBM J. Research and Development, 17(6):525–532, 1973.

[17] C. H. Bennett. The thermodynamics of computation, a review. International Journal of Theoretical Physics, 21(12):905–940, 1982.

[18] C. H. Bennett. Notes on the history of reversible computation. IBM J. Research and Development, 32(1):16–23, January 1988. Reprinted in [82], ch. 4, pp. 281– 288.

[19] C. H. Bennett. Time/space trade-offs for reversible computation. SIAM J. Computing, 18(4):766–776, 1989.

[20] Ethan Bernstein and Umesh V. Vazirani. Quantum complexity theory. In 25th Association for Computing Machinery Symposium on the Theory of Computing, pages 11–20, 1993.

[21] A. Berthiaume and Gilles Brassard. Oracle quantum computing. In Proceedings of the Workshop on Physics of Computation: PhysComp ’92, pages 195–199, Los Alamitos, CA, 1992. Institute of Electrical and Electronic Engineers Computer Society Press. Also to appear in Journal of Modern Optics.

[22] A. Berthiaume and Gilles Brassard. The quantum challenge to structural complexity theory. In Proceedings of the Seventh Annual Structure in Complexity Theory Conference, pages 132–137, Los Alamitos, CA, 1992. Institute of Electrical and Electronic Engineers Computer Society Press.

[23] Gianfranco Bilardi and Franco Preparata. Horizons of parallel computation. Technical Report CS-93-20, Brown University, May 1993. Also avail-

able on the web at http://www.cs.brown.edu/publications/techreports/
reports/CS-93-20.html.

[24] Bob Boothe. Algorithms for bidirectional debugging. Technical Report USM/CS-98-2-23, Computer Science Department, University of Southern Maine, 96 Falmouth St, Portland ME 04104-9300, February 1998. Author’s

email: boothe@cs.usm.maine.edu.

[25] A. R. Calderbank and Peter W. Shor. Good quantum error-correcting codes exist. Los Alamos Physics Preprint Archive, http://xxx.lanl.gov/abs/quantph/9512032, December 1995.

[26] C. S. Calude, J. Casti, and M. J. Dinneen, editors. Unconventional Models of Computation. Springer, 1998. (Proc. of the First International Conference on Unconventional Models of Computation (UMC’98), held at the University of Auckland, January 5–9, 1998).

[27] Ruknet Cezzar. Reversible computer apparatus and methods of constructing and utilizing same. U.S. Patent #5,469,550, 21 November 1995.

[28] Anantha P. Chandrakasan and Robert W. Brodersen. Low Power Digital CMOS Design. Kluwer Academic Publishers, 1995.

[29] Isaac L. Chuang and Raymond Laflamme. Quantum error correction by coding.

Los Alamos Physics Preprint Archive, http://xxx.lanl.gov/abs/quant-ph/

9511003, November 1995.

[30] Isaac L. Chuang and Yoshihisa Yamamoto. A simple quantum computer. Physical Review A, 52:3489–3496, 1995. Preprint at Los Alamos Physics Preprint

Archive, http://xxx.lanl.gov/abs/quant-ph/9505011.

[31] J. I. Cirac and P. Zoller. Quantum computations with cold trapped ions. Physical Review Letters, 74:4091–4094, 1995.

[32] Don Coppersmith and Edna Grossman. Generators for certain alternating groups with applications to cryptography. Society for Industrial and Applied Mathematics J. Appl. Math., 29(4):624–627, December 1975.

[33] David G. Cory, Amr F. Fahmy, and Timothy F. Havel. Ensemble quantum computing by nuclear magnetic resonance spectroscopy. Proceedings of the National Academy of Sciences of the United States of America, 94(5):1634–1639, 4 March

1997. See ftp://deas.ftp.harvard.edu/techreports/tr-10-96.ps.gz.

[34] David G. Cory, M. D. Price, and T. F. Havel. Nuclear magnetic resonance spectroscopy: An experimentally accessible paradigm for quantum computing. Physica D, 120(1–2):82–101, 1 September 1998.

[35] B.C. Crandall and James Lewis, editors. Nanotechnology: Research and Perspectives. MIT Press, 1992.

[36] Pierluigi Crescenzi and Christos H. Papadimitriou. Reversible simulation of space-bounded computation. Theoretical Computer Science, 143:159–165, 1995.

[37] Defense Advanced Research Projects Agency (DARPA). Scalable computing

systems. http://www.darpa.mil/ito/research/scalable.

[38] J. S. Denker, S. C. Avery, A. G. Dickinson, A. Kramer, and T. R. Wik. Adiabatic computing with the 2N-2N2D logic family. In International Workshop on Low Power Design, pages 183–187, 1994.

[39] David Deutsch and Richard Jozsa. Rapid solution of problems by quantum computation. Proceedings of the Royal Society of London Ser. A, A439:553– 558, 1992.

[40] Alexander G. Dickinson. Adiabatic logic. U.S. Patent #5,521,538, 28 May 1996. Assigned to AT&T.

[41] David P. DiVincenzo. Two-bit gates are universal for quantum computation. Physical Review A, 51(2):1015–1022, February 1995. Also at Los Alamos Physics

Preprint Archive, http://xxx.lanl.gov/abs/cond-mat/9407022.

[42] M. Dorojevets, P. Bunyk, D. Zinoviev, and K. Likharev. COOL-0: an RSFQ subsystem design for petaflops computing. IEEE Trans. on Appl. Supercond., jun 1999.

[43] K. Eric Drexler. Nanosystems: Molecular Machinery, Manufacturing, and Computation. John Wiley & Sons, Inc., 1992. http://nano.xerox.com/nanotech/

nanosystems.html.

[44] François Englert. On the black hole unitarity issue. Los Alamos e-print http:

//xxx.lanl.gov/abs/hep-th/9705115, 15 May 1997.

[45] T. Pellizzari et al. Decoherence, continuous observation and quantum computing: a cavity QED model. Preprint, June 1995. University of Innsbruck.

[46] Hugh Everett, III. The theory of the universal wave function. In Bryce S. De-Witt and Neill Graham, editors, The Many-Worlds Interpretation of Quantum Mechanics, pages 3–139. Princeton University Press, 1973.

[47] William Feller. An Introduction to Probability Theory and Its Applications. John Wiley & Sons, Inc., 1950.

[48] Richard Feynman. Quantum mechanical computers. Optics News, 11, 1985. Also in Foundations of Physics, 16(6):507–531, 1986.

[49] David J. Frank. CMOS toggle flip-flop using adiabatic switching. U.S. Patent #5,517,145, 4 May 1996. Assigned to IBM.

[50] David J. Frank. Static combinatortial logic circuits for reversible computation. U.S. Patent #5,493,240, 20 February 1996. Assigned to IBM.

[51] David J. Frank and Paul M. Solomon. Energy conserving clock pulse generating circuits. U.S. Patent #5,506,520, 9 April 1996. Assigned to IBM.

[52] Michael Frank. Time-symmetric control-flow instructions for less garbage in reversible programs. MIT Reversible Computing Project Internal Memo #M1,

February 1996. http://www.ai.mit.edu/~mpf/rc/memos/M01_symmarch.-
html.

[53] Michael Frank and Scott Rixner. Tick: A simple reversible processor (6.371 project report). Online term paper, May 1996. http://www.ai.mit.edu/~mpf/

rc/tick-report.ps.

[54] Michael P. Frank. Low-energy computing for implantable medical devices. Talk presented to the MIT Clinical Decision Making Group, http://www.ai.mit.-

edu/~mpf/rc/MedG_talk/webpage.html, 21 February 1996.

[55] Michael P. Frank. Modifications to PISA architecture to support guaranteed reveribility and other features. MIT Reversible Computing Project Internal

Memo #M7, February 1997. http://www.ai.mit.edu/~mpf/rc/memos/M07/
M07_revarch.html.

[56] Michael P. Frank. Voltage scaling and limits to energy efficiency for CMOSbased SCRL. MIT Reversible Computing Project Internal Memo #M3,

April 1997. Unfinished draft at http://www.ai.mit.edu/~mpf/rc/memos/
M03_scrllimits.html.

[57] Michael P. Frank and M. Josephine Ammer. Separations of reversible and irreversible space-time complexity classes. Extended abstract submitted to FOCS-

97. http://www.ai.mit.edu/~mpf/rc/memos/M06_oracle.html, 1997.

[58] Michael P. Frank and Thomas F. Knight, Jr. Ultimate theoretical models of nanocomputers. Nanotechnology, 9(3):162–176, 1998. Presented at the Fifth Foresight Conference on Molecular Nanotechnology, Palo Alto, CA, November

1997. http://www.ai.mit.edu/~mpf/Nano97/paper.html.

[59] Michael P. Frank, Thomas F. Knight, Jr., and Norman H. Margolus. Reversibility in optimally scalable computer architectures. In Calude et al. [26], pages

165–182. http://www.ai.mit.edu/~mpf/rc/scaling_paper/scaling.html.

[60] Michael P. Frank, Carlin Vieri, M. Josephine Ammer, Nicole Love, Norman H. Margolus, and Thomas F. Knight, Jr. A scalable reversible computer in silicon. In Calude et al. [26], pages 183–200. http://www.ai.mit.edu/~mpf/rc/

flattop/ft.html.

[61] E. F. Fredkin and T. Toffoli. Design principles for achieving high-performance submicron digital technologies. DARPA Proposal, November 1978.

[62] E. F. Fredkin and T. Toffoli. Conservative logic. International Journal of Theoretical Physics, 21(3/4):219–253, 1982.

[63] Ed Fredkin. Feynman, Barton and the reversible Schrödinger difference equation. In Hey [68], chapter 20, pages 337–348.

[64] Neil A. Gershenfeld and Isaac L. Chuang. Bulk spin resonance quantum computation. Science, 275(350), jan 1997.

[65] J. Storrs Hall. An electroid switching model for reversible computer architectures. In PhysComp92 [109], pages 237–247.

[66] J. Storrs Hall. An electroid switching model for reversible computer architectures. In Proc. ICCI ’92, 4th Int’l Conf. on Computing and Information, 1992.

[67] J. Storrs Hall. A reversible instruction set architecture and algorithms. In PhysComp94 [110], pages 128–134.

[68] Anthony J. G. Hey, editor. Feynman and Computation: Exploring the Limits of Computers. Perseus Books, Reading, MA, 1999.

[69] R. T. Hinman and M. F. Schlecht. Recovered energy logic: A single clock AC logic. In International Workshop on Low Power Design, pages 153–158, 1994.

[70] David A. Hodges and Horace G. Jackson. Analysis and Design of Digital Integrated Circuits. McGraw-Hill, Inc., second edition, 1988.

[71] John E. Hopcroft and Jeffrey D. Ullman. Introduction to Automata Theory, Languages, and Computation. Addison-Wesley Series in Computer Science. Addison-Wesley, 1979.

[72] K. Huang. Statistical Mechanics. Wiley, 1963.

[73] E. Joos and A. Qadir. A quantum statistical upper bound on entropy. Il Nuovo Cimento, 107B(5):563–572, 1992.

[74] Thomas F. Knight, Jr. and Saed Younis. Charge recovery logic including split level logic. U.S. Patent #5,378,940, 3 January 1995. Assigned to MIT.

[75] Tom Knight. An architecture for mostly functional languages. In Patrick Henry Winston and Sarah Alexandra Shellard, editors, Artificial Intelligence at MIT: Expanding Frontiers, volume 1, chapter 19, pages 500–519. The MIT Press, Cambridge, Massachusetts, 1990.

[76] J. G. Koller and W. C. Athas. Adiabatic switching, low energy computing, and the physics of storing and erasing information. In PhysComp92 [109], pages 267–270.

[77] A. Kramer, J. S. Denker, S. C. Avery, A. G. Dickinson, and T. R. Wik. Adiabatic computing with the 2N-2N2D logic family. In 1994 Symp. VLSI Circ.: Digest of Tech. Papers. Institute of Electrical and Electronic Engineers Press, June 1994.

[78] A. Kramer, J. S. Denker, B. Flower, and J. Moroney. 2nd order adiabatic computation with 2N-2P and 2N-2N2P logic circuits. In Intl. Symposium on Low Power Devices, pages 191–196, Dana Point, CA, 1995. ACM.

[79] Rolf Landauer. Irreversibility and heat generation in the computing process. IBM J. Research and Development, 5:183–191, 1961. Reprinted in [82], ch. 4, pp. 188–196.

[80] Klaus-Jörn Lange, Pierre McKenzie, and Alain Tapp. Reversible space equals deterministic space. In Proc. 12th Annual IEEE Conf. on Computational Complexity (CCC ’97), pages 45–50, June 1997. http://www.iro.umontreal.ca/

~tappa/Publications/LMT'97_abstract.html.

[81] Y. Lecerf. Machines de Turing réversibles. Insolubilité récursive en n ∈ N de l’équation u = θ^{n}, où θ est un ≪ isomorphisme de codes ≫ [Reversible Turing machines. Recursive insolubility in n ∈ N of the equation u = θ^{n}, where θ is an “isomorphism of codes”]. Comptes Rendus Hebdomadaires des Séances de L’académie des Sciences [Weekly Proceedings of the Academy of Science], 257:2597–2600, October 28, 1963. Unauthorized English translation at http:/

/www.ai.mit.edu/~mpf/rc/Lecerf/lecerf.html.

[82] Harvey S. Leff and Andrew F. Rex, editors. Maxwell’s demon: entropy, information, computing. Princeton series in physics. Princeton University Press, Princeton, NJ, 1990. May be ordered through http://www.ioppublishing.-

com/Books/Catalogue/020/__26/0750300566.

[83] F. Thomson Leighton. Introduction to Parallel Algorithms and Architectures: Arrays · Trees · Hypercubes. Morgan Kaufmann Publishers, San Mateo, California, 1992.

[84] Robert Y. Levine and Alan T. Sherman. A note on Bennett’s time-space tradeoff for reversible computation. SIAM J. Computing, 19(4):673–677, 1990.

[85] Ming Li and Paul Vitányi. An Introduction to Kolmogorov Complexity and Its Applications. Graduate Texts in Computer Science. Springer-Verlag, 2nd edition, 1997.

[86] Ming Li and Paul M. B. Vitányi. Reversibility and adiabatic computation: trading time and space for energy. Proceedings of the Royal Society of London Ser. A, 452:1–21, 1996.

[87] Ming Li and Paul M. B. Vitányi. Reversible simulation of irreversible computation. In Proc. 11th IEEE Conference on Computational Complexity, Philadelphia, Pennsylvania, May 24–27, 1996.

[88] K. K. Likharev. Classical and quantum limitations on energy consumption in computation. International Journal of Theoretical Physics, 21(3/4):311–326, [89] K. K. Likharev. Rapid single-flux-quantum logic. http://pavel.physics.-

sunysb.edu/RSFQ/Research/WhatIs/rsfqre2m.html, 1992.

[90] Konstantin K. Likharev. Rapid single flux quantum (RSFQ) superconduc-

tor electronics. See http://pavel.physics.sunysb.edu/RSFQ/Research/
WhatIs/rsfqwte1.html, may 1996.

[91] Konstantin K. Likharev and Alexander N. Korotkov. “Single-electron parametron”: Reversible computation in a discrete-state system. Science, 273:763–765, 9 August 1996.

[92] Chris Lutz. Janus: A time-reversible language. Letter from Chris Lutz to Rolf Landauer. Unauthorized reproduction at http://www.ai.mit.edu/~mpf/rc/ janus.html, 1 April 1986.

[93] N. H. Margolus. Physics and Computation. PhD thesis, Massachusetts Institute of Technology, 1988.

[94] Norman Margolus. Physics-like models of computation. Physica D, 10:81–95, 1984.

[95] Norman Margolus. Crystalline computation. In Hey [68], chapter 18, pages 267–305.

[96] Norman Margolus and Lev B. Levitin. The maximum speed of dynamical evolution. In Toffoli et al. [137], pages 208–211. Available through http://www.-

interjournal.org. Revised version available at ftp://im.lcs.mit.edu/poc/
margolus/speed.of.dynamics.ps.Z.

[97] Norman Margolus, Tommaso Toffoli, and Gérard Vichniac. Cellular-automata supercomputers for fluid dynamics modeling. Phys. Rev. Lett., 56(16):1694– 1696, 21 April 1986.

[98] Vladimir S. Mashkevich. Conservative model of black hole (sic) and lifting of the information loss paradox. Los Alamos e-print http://xxx.lanl.gov/abs/

gr-qc/9707055, 27 July 1997.

[99] Carver A. Mead. Scaling of MOS technology to submicrometer feature sizes. Journal of VLSI Signal Processing, 8:9–25, 1994. Reprinted as chapter 9 of [68].

[100] Ralph C. Merkle. Towards practical reversible logic. In PhysComp92 [109], pages 227–228.

[101] Ralph C. Merkle. Reversible electronic logic using switches. Nanotechnology, 4:21–40, 1993.

[102] Ralph C. Merkle. Two types of mechanical reversible logic. Nanotechnology, 4:114–131, 1993.

[103] Ralph C. Merkle. Reversible charge transfer and logic utilizing them. U.S. Patent #5,357,548, 18 October 1994. Assigned to Xerox.

[104] Ralph C. Merkle and K. Eric Drexler. Helical logic. Nanotechnology, 7(4):325–

339, 1996. Available through http://www.ioppublishing.com.

[105] C. D. Motchenbacher and J. A. Connelly. Low-Noise Electronic System Design. John Wiley & Sons, Inc., 1993.

[106] Robert C. Myers. Pure states don’t wear black. Los Alamos e-print http://

xxx.lanl.gov/abs/gr-qc/9705065, May 1997.

[107] William J. Ooms and Jerald A. Hallmark. Complementary logic recovered energy circuit. U.S. Patent #5,426,382, 20 June 1995. Assigned to Motorola.

[108] Christos H. Papadimitriou. Computational Complexity. Addison-Wesley, 1994.

[109] PhysComp ’92: Proceedings of the Workshop on Physics and Computation, October 2–4, 1992, Dallas, Texas, Los Alamitos, CA, 1992. IEEE Computer Society Press.

[110] PhysComp ’94: Proceedings of the Workshop on Physics and Computation, November 17–20, 1994, Dallas, Texas, Los Alamitos, CA, 1994. IEEE Computer Society Press.

[111] J. E. Pin. On the languages accepted by finite reversible automata. In Thomas Ottman, editor, Automata, Languages and Programming, Proc. 14th Int’l Colloq. (ICALP), volume 267 of Lecture Notes in Computer Science, pages 237–249. Springer-Verlag, 1987.

[112] M. Planck. The Theory of Heat Radiation. Dover, New York, 1991.

[113] John Preskill. Do black holes destroy information? Los Alamos e-print http:

//xxx.lanl.gov/abs/hep-th/9209058, September 1992.

[114] J.M. Rabaey. Digital Integrated Circuits: A Design Perspective. Prentice-Hall,

1996. http://infopad.EECS.Berkeley.EDU/~icdesign.

[115] A. L. Ressler. The design of a conservative logic computer and a graphical editor simulator. Master’s thesis, MIT Artificial Intelligence Laboratory, 1981.

[116] Hartley Rogers, Jr. Theory of Recursive Functions and Effective Computability. The MIT Press, first mit press paperback edition edition, 1987. Original edition published by McGraw-Hill Book Company, 1967.

[117] Martin F. Schlecht and Roderick T. Hinman. Recovered energy logic circuits. U.S. Patent #5,396,527, 7 March 1995. Assigned to MIT.

[118] Francis W. Sears, Mark W. Zemansky, and Hugh D. Young. University Physics. Addison-Wesley Series in Physics. Addison-Wesley Publishing Company, Reading, Massachusetts, sixth edition, February 1984.

[119] Charles L. Seitz, Alexander H. Frey, Sven Mattisson, Steve D. Rabin, Don A. Speck, and Jan L. A. van de Snepscheut. Hot-clock nMOS. In Henry Fuchs, editor, 1985 Chapel Hill Conference on Very Large Scale Integration, pages 1– 17. Computer Science Press, 1985.

[120] Semiconductor Industry Association. The National Technology Roadmap for Semiconductors: Technology Needs. SEMATECH, Inc., 1997 edition, 1997.

http://notes.sematech.org/ntrs/Rdmpmem.nsf.

[121] Peter W. Shor. Algorithms for quantum computation: Discrete log and factoring. In Proceedings of the 35th Annual Symposium on Foundations of Computer Science, pages 124–134. Institute of Electrical and Electronic Engineers Com-

puter Society Press, November 1994. ftp://netlib.att.com/netlib/att/
math/shor/quantum.algorithms.ps.Z.

[122] Hava (Eve) Tova Siegelmann. Foundations of Recurrent Neural Networks. PhD thesis, Rutgers University, 1993.

[123] David R. Simon. On the power of quantum computation. In Proceedings of the 35th Annual Symposium on Foundations of Computer Science, pages 116– 123, Los Alamitos, CA, 1994. Institute of Electrical and Electronic Engineers

Computer Society Press. Preprint at http://vesta.physics.ucla.edu/cgi-
bin/uncompress_ps_cgi?simon94.ps. Improves on Bernstein & Vazirani ’93

[20]. Shor ’94 [121] was inspired by this.

[124] Tycho Sleator and Harald Weinfurter. Realizable universal quantum logic gates. Physical Review Letters, 74:4087–4090, 1995. Proposes a universal gate and a Cavity QED implementation.

[125] Warren D. Smith. Church’s thesis meets the N-body problem. Technical Report TM 93-105-3-0058-6, NECI, September 1993. http://www.neci.nj.nec.com/

homepages/wds/church.ps.

[126] Warren D. Smith. Fundamental physical limits on computation. Technical

report, NECI, May 1995. http://www.neci.nj.nec.com/homepages/wds/
fundphys.ps.

[127] Paul M. Solomon and David J. Frank. Power measurements of adiabatic circuits by thermoelectric technique. In Symposium on Low Power Electronics, pages 18–19, 1995.

[128] Dinesh Somasekhar, Yibin Ye, and Kaushik Roy. An energy recovery static RAM memory core. In Symposium on Low Power Electronics, pages 62–63, 1995.

[129] Andrew Steane. Multiple particle interference and quantum error correction. Proceedings of the Royal Society of London Ser. A, 1996. (Submitted.) Preprint available at Los Alamos Physics Preprint Archive, http://xxx.lanl.gov/abs/ quant-ph/9601029. Introduces a parity encoding for quantum error correction.

[130] L. “J.” Svensson and J. G. Koller. Adiabatic charging without inductors. In Proc. Int’l Workshop on Low-Power Design, pages 159–164, apr 1994.

[131] Lars Svensson. Adiabatic switching. In Low Power Digital CMOS Design [28], chapter 6, pages 181–218.

[132] W. G. Teich, K. Obermayer, and G. Mahler. Structural basis of multistationary quantum systems ii: Effective few-particle dynamics. Physical Review B, 37:8111–8121, 1988.

[133] Steven D. Thomas. Precharged adiabatic pipelined logic. U.S. Patent #5,602,497, 11 February 1997.

[134] Tommaso Toffoli. Computation and construction universality of reversible cellular automata. J. Computer and System Sciences, 15:213–231, 1977.

[135] Tommaso Toffoli. Reversible computing. Technical memo MIT/LCS/TM-151, MIT Lab for Computer Science, February 1980. Out of print; available from NTIS. Abridged version available as [136].

[136] Tommaso Toffoli. Reversible computing. In J. W. de Bakker and J. van Leeuwen, editors, Automata, Languages and Programming (Seventh Colloquium, Noordwijkerhout, the Netherlands, July 14–18, 1980), volume 85 of Lecture Notes in Computer Science, pages 632–644. Springer-Verlag, 1980. Abridged version of [135].

[137] Tommaso Toffoli, Michael Biafore, and João Leão, editors. PhysComp96 (Proceedings of the Fourth Workshop of Physics and Computation, Boston University, 22–24 November 1996). New England Complex Systems Institute, 1996. Copies may be ordered from PhysComp96, 44 Cummington St., Boston MA 02215 (PhysComp96@pm.bu.edu). Individual papers are available through http:

//www.interjournal.org.

[138] Tommaso Toffoli and Norman Margolus. Cellular Automata Machines: A New Environment for Modeling. MIT Press, 1987.

[139] Kai-Yap Toh, Ping-Keung Ko, and Robert G. Meyer. An engineering model for short-channel MOS devices. IEEE Journal of Solid-State Circuits, 23(4):950– 958, August 1988.

[140] D. B. Tuckerman and R. F. W. Pease. High-performance heat sinking for VLSI. IEEE Electron Device Letters, EDL-2(5):126–129, May 1981.

[141] David Bazeley Tuckerman. Heat-Transfer Microstructures for Integrated Circuits. PhD thesis, Stanford University, February 1984.

[142] Q. A. Turchette, C. J. Hood, W. Lange, H. Mabuchi, and H. Jeffrey Kimble. Measurement of conditional phase shifts for quantum logic. Submitted to

Physical Review Letters. Abstract at http://www.cco.caltech.edu/~hood/
QO/Abstracts/Turc95b.html.

[143] A. M. Turing. On computable numbers, with an application to the entscheidungsproblem. Proc. London Math. Society, 2(42):230–265, 1936. Also no. 43, pp. 544–546, 1937.

[144] N. Tzartzanis and W. Athas. Clock-powered CMOS: A hybrid adiabatic logic style for power-efficient computing. In 20th Anniversary Conference on Advanced Research in VLSI, pages 137–151. IEEE Computer Society Press, March

1999. Slides at http://www.isi.edu/acmos/presentations/99-03.ARVLSI.-
pdf.

[145] Nestoras Tzartzanis and William C. Athas. Energy recovery for the design of high-speed, low power static RAMs. In International Symposium on Low Power Electronics and Design, pages 55–60, 1996.

[146] David Ungar, Henry Lieberman, and Christopher Fry. Debugging and the experience of immediacy. Communications of the ACM, 40(4):38–43, April 1997.

[147] P. van Emde Boas. Machine models and simulations. In J. van Leeuwen, editor, Handbook of Theoretical Computer Science, volume A, pages 1–66. Elsevier, Amsterdam, 1990.

[148] A. Vergis, K. Steiglitz, and B. Dickinson. The complexity of analog computation. Math. and Computers in Simulation, 28:91–113, 1986. .

[149] Anastasios Vergis, Kenneth Steiglitz, and Bradley Dickinson. The complexity of analog computation. Mathematics and Computers in Simulation, 28:91–113, 1986.

[150] Carlin Vieri. Reversible Computer Engineering and Architecture. PhD thesis, Massachusetts Institute of Technology, 1999.

[151] Carlin J. Vieri. Pendulum: A reversible computer architecture. Master’s thesis, MIT Artificial Intelligence Laboratory, 1995.

[152] Paul M. B. Vitányi. Locality, communication and interconnect length in multicomputers. SIAM J. Computing, 17:659–672, 1988.

[153] J. von Neumann. Non-linear capacitance or inductance switching, amplifying, and memory organs. U.S. Patent #2,815,488, December 3, 1957. Assigned to IBM.

[154] John von Neumann. Theory of Self-Reproducing Automata. University of Illinois Press, 1966.

[155] L. Wang. On the classical limit of phase-space formulation of quantum mechanics: entropy. J. Math. Phys., 27:483–487, 1986.

[156] Steve Ward, John Nguyen, and John Pezaris. 3D-3N meshes. MIT LCS NuMesh project internal memo, sep 1991. http://www.cag.lcs.mit.edu/

numesh/papers/memos/3d3n.html.

[157] Boyd G. Watkins. A low-power multiphase circuit technique. IEEE Journal of Solid-State Circuits, pages 213–220, December 1967.

[158] Neil H. E. Weste and Kamran Eshraghian. Principles of CMOS VLSI Design: A Systems Perspective. Addison-Wesley, second edition, 1993.

[159] R. L. Wigington. A new concept in computing. Proceedings of the IRE, 47:516– 523, April 1961.

[160] Yibin Ye and Kaushik Roy. Energy recovery circuits using reversible and partially reversible logic. IEEE Transactions on Circuits and Systems—I: Fundamental Theory and Applications, 43(9):769–778, sep 1996.

[161] S. G. Younis. Asymptotically Zero Energy Computing Using Split-Level Charge Recovery Logic. PhD thesis, MIT Artificial Intelligence Laboratory, 1994.

[162] S. G. Younis and T. F. Knight, Jr. Practical implementation of charge recovering asymptotically zero power CMOS. In Proc. 1993 Symp. on Integrated Systems, pages 234–250. MIT Press, 1993.

[163] S. G. Younis and T. F. Knight, Jr. Asymptotically zero energy split-level charge recovery logic. In International Workshop on Low Power Design, pages 177–182, 1994.

[164] S. G. Younis and T. F. Knight, Jr. Harmonic resonant rail drivers for adiabatic logic. In Proceedings of the 1995 Symposium on Advanced Research in VLSI. MIT Press, 1995.

[165] D. Zinoviev. Design issues in ultra-fast ultra-low-power superconductor batcherbanyan switching fabric based on RSFQ logic/memory family. Applied Superconductivity, 5(7–12):235–239, aug 1998.

That’s all, folks!

406 BIBLIOGRAPHY